OpenStudy (anonymous):

some one please solve g+4-3g=1+g

OpenStudy (anonymous):

g+4-3g=1+g combine like terms -2g+4=1+g -3g+4=1 -3g=-3 g=1

OpenStudy (anonymous):

thanks :)